a. Determine the x and y components of reaction at C using scalar notation.

Express your answers using three significant figures separated by a comma.

b. Determine the x and z components of reaction at B using scalar notation.

Express your answers using three significant figures separated by a comma.

c.Determine the x and z components of reaction at A using scalar notation.

Express your answers using three significant figures separated by a comma.
Answer #1

Fcy = 590 N (Ans)

Taking moment of the system,

-0.6 Fbz - 370 x 1.2 - Fcy x 0.4 = 0

Fbz = - 1133.33 N (Ans)

Faz = - 370 + 1133.33 = 763.33 N (Ans)

Taking moment about u axis

-370 x 0.6 - Fcx x 0.4 = 0

Fcx = - 555 N (Ans)

-Fbx x 0.6 + Fcy x 0.6 - Fcx x 1.2 = 0

Fbx = 520 N (Ans)

Fax = - 520 + 555 = 35 N (Ans)